# Training Neural Networks on Data Sources with Unknown Reliability

### v2

@@ A.4 Alternative interpretation
As discussed in Appendix A.2, the model evidence, $p(\mathcal{D}|\mathcal{M})$ ($\mathcal{D}$: the dataset and $\mathcal{M}$: the model) is the probability that a dataset is generated by a given model (by marginalisation of the model parameters). In our case, this could be used to calculate the probability that data from each source, $s$ is generated by the given model, namely: $p(\mathcal{D}_{s}|\mathcal{M})$ where $\mathcal{D}_{s}$ is the data generated by a source $s$. A model designed for non-noisy data should have a high $p(\mathcal{D}_{s}|\mathcal{M})$ when $\mathcal{D}_{s}$ has a low noise level, and a low $p(\mathcal{D}_{s}|\mathcal{M})$ when $\mathcal{D}_{s}$ has a high noise level. However, $\log(p(\mathcal{D}_{s}|\mathcal{M}))$ is hard to compute for neural networks, and would require an approximation. It can be shown that $p(\mathcal{D}|\mathcal{M})$ can be approximated by the logsumexp of the log-likelihood over a training trajectory with noisy gradients, which approximates the posterior samples with SGLD or SGHMC [42, 43]. Within our work, when computing if a given source is noisy or non-noisy, the mean loss for this source is calculated over a training trajectory of length equal to the history length, $H$ (given by $\mu_{s}$ in Algorithm 1). This is then compared to a weighted mean of the loss over a training trajectory of length $H$ of all sources excluding the source under scrutiny (given by $\mu_{s^{\prime}}$ in Algorithm 1). This weighted mean is related to $p(\mathcal{D}_{R}|\mathcal{M})$ where $\mathcal{D}_{R}$ represents the data from all non-noisy sources, since the weights allow us to reduce the influence of noisy sources. Therefore, our method is related to one in which the model evidence is calculated for each source at each step, and where the sources with a relative model evidence higher than some threshold (defined in our work through the leniency, $\lambda$) have larger influence on the model parameters when updating the weights.
A.5 Details of the varied parameter experiments
To build an intuition for the parameters introduced in defining LAP, we run some synthetic examples to understand their effects on $1-d_{s}$ (in Equation 6). The results of this are presented in Figure 2.
Dataset. The synthetic data is created using Scikit-Learn’s make_moons  [^3] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_moons.html function, which produced $10,000$ synthetic observations with $2$ features. Each observation is assigned a label based on which “moon" the observation corresponded to. Then, to produce synthetic sources and noise levels for each source, we randomly assign each data point a source number from 00 to $4$, so that we have $5$ sources in total. Data points for each source were then made noisy by randomly flipping labels such that each source had a noise level of $0.0$, $0.025$, $0.05$, $0.25$, or $1.0$.
@@ A.5 Details of the varied parameter experiments
These experiments take approximately $5$ minutes to complete using the compute described in Appendix A.1.
A.6 Further information on the datasets
@@
  • CIFAR-10 and CIFAR-100 [30] are datasets made up of $60000$ RGB images of size $32\times 32$ divided into 10 and 100 classes, with $6000$ and $600$ images per class respectively. The task in this dataset is to predict the correct class of a given image.
  • F-MNIST [31] is made up of $70000$ greyscale images of clothes of size $28\times 28$ divided into 10 classes with $7000$ images per class, which is flattened into observations with 784 features. The task of this dataset is to predict the correct class of a given image.
@@
  • CIFAR-10N [33] is a dataset made up of all examples from CIFAR-10, but with human labelling categorised into different levels of quality. For our case, to allow for the most flexibility in experiment design, we utilised the worst of the human labels.
  • IMDB [36] is a natural language dataset with a sentiment analysis task. The goal is to correctly classify a movie review as either positive or negative based on its text. It contains $25000$ reviews in the training set, and $25000$ reviews in the testing set, split equally between positive and negative sentiment.
  • Tiny-Imagenet [32] contains $110000$ RGB images of size $64\times 64$ from $200$ classes, with the goal of correctly classifying an image into the given class. There are $100000$ training images and $10000$ images in the test set.
@@
  • California Housing [37] contains $20640$ samples of house values in California districts, with the goal to predict the median house value from the U.S census data for that region. We randomly split the data into training and testing with an $8:2$ ratio on each repeat.
Dataset licenses. CIFAR-10, CIFAR-100, and F-MNIST are managed under the MIT License, whilst PTB-XL is made available with the Creative Commons Attribution 4.0 International Public License  [^4] https://creativecommons.org/licenses/by/4.0/., and CIFAR-10N is available with the Creative Commons Attribution-Non Commercial 4.0 International Public License  [^5] https://creativecommons.org/licenses/by-nc/4.0/.. IMDB [^6] https://ai.stanford.edu/~amaas/data/sentiment/ and California Housing [^7] https://www.dcc.fc.up.pt/~ltorgo/Regression/cal_housing.html are made publicly available, and Tiny-Imagenet and Imagenet are available after agreeing to the terms of access  [^8] https://www.image-net.org/download.
Dataset availibility. As mentioned in Appendix A.1, all datasets tested are publicly available and easily accessible. Additionally, within the supplementary code we use the default Pytorch Dataset objects for CIFAR-10, CIFAR-100, and F-MNIST and provide Pytorch Dataset objects that will automatically download, unzip, and load the data for PTB-XL, CIFAR-10N, IMDB, and California Housing. We additionally provide code to load Tiny-Imagenet and Imagenet from a local directory, since it requires an agreement before accessing. This makes reproduction of the work presented simple to perform.
A.7 Experiments in detail
@@
Within our experiments, we tested $4$ model architectures, described below. Due to limitations in compute and to allow for our work to be easily reproducible, we use different levels of model capacity which additionally illustrates that our method is applicable in varied settings. All models are also available in the supplementary code, implemented in Pytorch.
@@
  • ResNet 2D: This model is designed for image classification and is exactly the ResNet 18 or Resnet 50 architecture presented in He et al. [38]. For our implementation, we used the model source code available with Pytorch  [^9] https://pytorch.org/vision/main/models/resnet.html.
  • LSTM: The natural language model contains an embedding layer, which maps tokens to vectors of size $256$, a predefined LSTM module from Pytorch  [^10] https://pytorch.org/docs/stable/generated/torch.nn.LSTM.html with $2$ layers and a hidden size of $512$, a dropout layer with a probability of $0.25$, and a fully connected layer mapping the output from the LSTM module to $2$ classes.
@@
  • CIFAR-10 and CIFAR-100: All training data is randomly split into $10$ sources, with $4$ and $2$ sources chosen to be noisy for CIFAR-10 and CIFAR-100 respectively. In this experiment, these noisy sources are chosen to be $100\%$ noisy so that we can get an understanding of how our method performs on data containing highly noisy sources. Noise is introduced based on the description given in Section 3. The Low capacity CNN described above is trained on this data and tasked with predicting the image class using cross entropy loss. The model is trained for $25$ epochs in both cases, with the Adam optimiser [44] and a learning rate of $0.001$ on batches of size $128$. When training with LAP, we use $H=25$, $\delta=1.0$, and $\lambda=0.8$ chosen using the validation data (which is made noisy using the same procedure as the training data) after trained on data with batch label flip noise. When training on CIFAR-100 we found that using a warm-up of $100$ steps improved performance.
  • F-MNIST: All training data is randomly split into $10$ sources, with $6$ chosen to be $100\%$ noisy. The low capacity MLP model is trained for $40$ epochs using the Adam optimiser, and a learning rate of $0.001$ on batches of size $200$. When training with LAP, we use $H=50$, $\delta=1.0$, and $\lambda=0.8$ chosen using the validation data (which is made noisy using the same procedure as the training data) after trained on data with batch label flip noise.
In our implementation, each data point contained an observation, the generating data source, and a target. In this experiment, to maintain consistency with ARFL, data points were assigned to mini-batches such that each batch contained data from a single source. The features and labels were passed to the model for training, and the sources were used to calculate the weighting to apply to the loss on each data point.
To allow fairness in the comparison with ARFL, global updates on this model were performed the same number of times as the number of epochs other models were trained for, with all clients being trained on data for a single epoch before each global update. All other parameters for the ARFL model are kept the same. When testing IDPA, the parameters were chosen as in Wang et al. [15] by using the default parameters in the implementation. Finally, when testing “Co-teaching", the parameters are chosen as in Han et al. [14] where possible, and scaled proportionally by the change in the number of epochs between their setting and our setting where they depended on the total number of epochs. The forgetting rate was set as the default given in the implementation code, $0.2$, since we do not assume access to the true noise rate.
@@
For all of the following experiments, to reduce computational cost, we fix the parameters of LAP to $H=25$, $\delta=1.0$, and $\lambda=0.8$. In practise, optimising these parameters should allow for improved performance when using LAP over those presented in this work.
Results in Table 2. To produce the results presented here, all training data is randomly split into $10$ sources, with $4$ chosen to be $100\%$ noisy. The high capacity CNN with contrastive learning presented in Li et al. [16] is trained for $25$ epochs, with all other parameters kept as in the original work. In addition to this model, a version (with the same parameters) is trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. These models were trained on batch sizes of $128$ using stochastic gradient descent with a learning rate of $0.02$, momentum of $0.9$, and weight decay $0.0005$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
@@
Results in Figure 3(a). This experiment allowed us to test the performance of using LAP on different numbers of sources with varied noise levels. Firstly, data is split into $12$ sources based on the clinician performing the labelling of the ECG recording. Then, for a given number of sources (increasing along the $x$ axis), the noisy sources have noise levels that are set at linear intervals between $25\%$ and $100\%$. These sources are made noisy following suggestions in Wong et al. [35] to simulate electromagnetic interference and using label flipping to simulate human error in labelling. On this data, we train the ResNet 1D model for $40$ epochs using the Adam optimiser with a learning rate of $0.001$ and a batch size of $64$. This model is trained with and without LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
@@
Results in Figure 3(b). Here, CIFAR-10 is used as the features and the labelling collected in Wei et al. [33] are used as the noisy labels to produce CIFAR-10N. The data is first randomly split into $10$ sources. Then, for a given number of sources (increasing along the $x$ axis), the noisy sources have noise levels that are set at linear intervals between $25\%$ and $100\%$, by replacing the true CIFAR-10 labels with the real noisy labels from Wei et al. [33]. The high capacity CNN with contrastive learning presented in Li et al. [16] is trained for $25$ epochs, with all other parameters kept as in the original work. As before, a version of this model (with the same parameters) is trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. These models are trained on batch sizes of $128$ using stochastic gradient descent with a learning rate of $0.02$, momentum of $0.9$, and weight decay $0.0005$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
@@
Results in Table 3. In this set of experiments, we wanted to test the use of LAP in a natural language setting. Firstly, we load the training and testing sets and randomly split the training set into $10$ sources uniformly. We then choose $4$ of the sources to be $100\%$ unreliable, and introduce noise through random labelling and randomly permuting the order of the text. We truncate or extend each review such that it contains exactly $256$ tokens. We use an LSTM to predict the sentiment of the movie reviews by training the model for $40$ epochs with a batch size of $128$, leaning rate of $0.001$ and the Adam optimiser [44]. A version of this model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources. Since IDPA and Co-teaching are not bench-marked on this dataset, for the baseline specific parameters we used the same values as given for CIFAR-10.
@@
A.8 Table 1 with multiple sources in batches
Table 4: Table 1 CIFAR-10 results, with multiple sources in each batch. Mean ± standard deviation of maximum test accuracy (%) of $5$ repeats of the baselines and LAP on synthetic data with different noisy types on CIFAR-10. There are $4$ noisy sources out of $10$ total sources. Unreliable sources are each 100% noisy. Note there are non-significant differences in results on “Original Data" between Standard and LAP training since here the data is non-noisy.
Model Types
Noise Type Standard IDPA Co-teaching LAP (Ours)
Original Data 69.46 ± 1.12 70.86 ± 0.32 67.73 ± 0.83 69.6 ± 1.08
Chunk Shuffle 65.92 ± 1.32 66.34 ± 0.99 63.55 ± 0.41 66.92 ± 0.56
Random Label 60.85 ± 1.67 57.93 ± 1.59 61.37 ± 1.01 66.5 ± 0.43
Batch Label Shuffle 62.23 ± 0.62 60.08 ± 0.13 63.9 ± 0.44 66.09 ± 0.65
Batch Label Flip 59.94 ± 2.48 61.02 ± 1.9 63.05 ± 2.43 64.48 ± 1.11
Added Noise 59.86 ± 1.38 61.19 ± 0.33 59.59 ± 1.37 63.58 ± 1.03
Replace With Noise 65.07 ± 0.63 65.85 ± 0.64 64.1 ± 1.13 64.89 ± 1.28
Since ARFL is designed for a federated learning setting, it is only able to access data from a single source at a time. As mentioned in Section 4, to maintain a consistent experimental setting, we therefore trained all methods in Table 1 on batches that contained only a single source. For completeness, and to highlight that the same conclusions still remain when batches contain multiple sources, we reproduce the results for CIFAR-10 but with multiple sources in each batch. These results are presented in Table 4. This table highlights the same pattern, but does additionally show that the noisy label baselines perform better than the standard training when batches contain multiple sources, likely because each batch has less variability in the number of noisy data points compared to the experiments in Table 1. Note in all experiments presented in this work, apart from Table 1, each batch contains multiple sources.

### v4

@@ A.6 Alternative interpretation
As discussed in Appendix A.2, the model evidence, $p(\mathcal{D}|\mathcal{M})$ ($\mathcal{D}$: the dataset and $\mathcal{M}$: the model) is the probability that a dataset is generated by a given model (by marginalisation of the model parameters). In our case, this could be used to calculate the probability that the data from each source $s$ is generated by the given model, that is, $p(\mathcal{D}_{s}|\mathcal{M})$ where $\mathcal{D}_{s}$ are the data generated by a source $s$. A model designed for non-noisy data should have a high $p(\mathcal{D}_{s}|\mathcal{M})$ when $\mathcal{D}_{s}$ has a low noise level and a low $p(\mathcal{D}_{s}|\mathcal{M})$ when $\mathcal{D}_{s}$ has a high noise level. However, $\log(p(\mathcal{D}_{s}|\mathcal{M}))$ is hard to compute for neural networks and would require an approximation. It can be shown that $p(\mathcal{D}|\mathcal{M})$ can be approximated by the logsumexp of the log-likelihood over a training trajectory with noisy gradients, which approximates posterior samples with SGLD or SGHMC [46, 47]. Within our work, when computing if a given source is noisy or non-noisy, the mean loss for this source is calculated over a training trajectory of length equal to the history length, $H$ (given by $\mu_{s}$ in Algorithm 1). This is then compared to a weighted mean of the loss over a training trajectory of length $H$ of all sources excluding the source under scrutiny (given by $\mu_{s^{\prime}}$ in Algorithm 1). This weighted mean is related to $p(\mathcal{D}_{R}|\mathcal{M})$ where $\mathcal{D}_{R}$ represents the data from all non-noisy sources, since the weights allow us to reduce the influence of noisy sources. Therefore, our method is related to one in which the model evidence is calculated for each source at each step, and where the sources with a relative model evidence higher than some threshold (defined in our work through leniency, $\lambda$) have a larger influence on the model parameters when updating the weights.
A.7 Details of the varied parameter experiments
To build an intuition for the parameters introduced in the definition of LAP, we run some synthetic examples to understand their effects on $1-d_{s}$ (in Equation 5). The results of this are presented in Figure 4.
Dataset. The synthetic data is created using Scikit-Learn’s make_moons  [^2] Scikit-Learn:make_moons function, which produced $10,000$ synthetic observations with $2$ features. Each observation is assigned a label based on which “moon” the observation corresponded to. Then, to produce synthetic sources and noise levels for each source, we randomly assign each data point a source number from 00 to $4$, so that we have $5$ sources in total. The data points for each source were then made noisy by randomly flipping labels such that each source had a noise level of $0.0$, $0.025$, $0.05$, $0.25$, or $1.0$.
@@ A.7 Details of the varied parameter experiments
These experiments take approximately $5$ minutes to complete when using the compute described in Appendix A.1.
A.8 Further information on the datasets
@@
  • CIFAR-10 and CIFAR-100 [32] are datasets made up of $60000$ RGB images of size $32\times 32$ divided into 10 and 100 classes, with $6000$ and $600$ images per class, respectively. The task in this dataset is to predict the correct class of a given image.
  • F-MNIST [33] is made up of $70000$ greyscale images of clothing of size $28\times 28$ divided into $10$ classes with $7000$ images per class, which are flattened into observations with $784$ features. The task of this dataset is to predict the correct class of a given image.
@@
  • CIFAR-10N [35] is a dataset made up of all CIFAR-10 examples, but with human labelling categorised into different levels of quality. For our case, to allow for the most flexibility in experiment design, we utilised the worst of the human labels.
  • GoEmotions [39] is a natural language dataset in which the task is to correctly classify the emotion of Reddit comments from a possible $28$ emotions. This dataset contains $171820$ text examples, annotated by $82$ raters, with each rater contributing somewhere between $1$ and $9320$ labels (mean: $1676$, standard deviation: $1477$). To produce the training and test set, we randomly split the examples in the 80:20 ratio.
  • IMDB [38] is a natural language dataset with a sentiment analysis task. The goal is to correctly classify a movie review as positive or negative based on its text. It contains $25000$ reviews in the training set and $25000$ reviews in the testing set, divided equally between positive and negative sentiment.
  • MNIST [48] contains $60000$ training and $10000$ testing images of handwritten digits in black and white with a resolution of $28$. The goal of this dataset is to correctly identify the digit drawn in the image (from 00 to $9$).
  • Tiny-Imagenet [34] contains $110000$ RGB images of size $64\times 64$ of $200$ classes, with the goal of correctly classifying an image into the given class. There are $100000$ training images and $10000$ images in the test set.
@@
  • California Housing [40] contains $20640$ samples of house values in California districts, with the aim of predicting the median house value from the US census data for that region. We randomly split the data into training and testing with an $8:2$ ratio on each repeat.
Dataset licenses. CIFAR-10, CIFAR-100, and F-MNIST are managed under the MIT License. PTB-XL is made available with the Creative Commons Attribution 4.0 International Public License  [^3] https://creativecommons.org/licenses/by/4.0/., CIFAR-10N is available with the Creative Commons Attribution-Non-Commercial 4.0 International Public License  [^4] https://creativecommons.org/licenses/by-nc/4.0/., GoEmotions is made available under the Apache-2.0 license  [^5] https://github.com/google-research/google-research., and MNIST is available under the the Creative Commons Attribution-Share-Alike 3.0 license  [^6] https://creativecommons.org/licenses/by-sa/3.0/. IMDB [^7] https://ai.stanford.edu/~amaas/data/sentiment/ and California Housing [^8] https://www.dcc.fc.up.pt/~ltorgo/Regression/cal_housing.html are made publicly available, and Tiny-Imagenet and Imagenet are available after agreeing to the access terms  [^9] https://www.image-net.org/download.
Dataset availibility. As mentioned in Appendix A.1, all datasets tested are publicly available and easily accessible. Additionally, within the supplementary code we use the default Pytorch Dataset objects for CIFAR-10, CIFAR-100, F-MNIST, and MNIST, and provide Pytorch Dataset objects that will automatically download, unzip, and load the data for PTB-XL, CIFAR-10N, GoEmotions, IMDB, and California Housing. We additionally provide code to load Tiny-Imagenet and Imagenet from a local directory since it requires an agreement before accessing. This makes reproduction of the work presented easy to perform.
A.9 Experiments in detail
@@
Within our experiments, we tested many model architectures, described below. Due to limitations in computation and to allow for our work to be easily reproducible, we use different levels of model capacity, which additionally illustrates that our method is applicable in varied settings. All models are also available in the supplementary code, implemented in Pytorch.
@@
  • ResNet 2D: This model is designed for image classification and is exactly the ResNet 20 architecture presented in He et al. [41] or the ResNet 18 or Resnet 50 implementation in available with Pytorch  [^10] Pytorch:ResNet.
  • Transformer Encoder: [43] This model is designed for the natural language based emotion prediction task given by the GoEmotions dataset. This model uses an embedding layer of size $256$, positional encoding, and $2$ transformer encoder layers (based on the implementation in Pytorch  [^11] Pytorch:TransformerEncoderLayer) with $4$ heads, and an embedding size of $256$. This is followed by a linear layer that maps the output from the transformer encoder to the $28$ emotion classes.
  • LSTM: [42] The natural language model contains an embedding layer, which maps tokens to vectors of size $256$, a predefined LSTM module from Pytorch  [^12] Pytorch:LSTM with $2$ layers and a hidden size of $512$, a dropout layer with a probability of $0.25$, and a fully connected layer mapping the output from the LSTM module to $2$ classes.
@@
  • CIFAR-10 and CIFAR-100: All training data is randomly split into $10$ sources, with $4$ and $2$ sources chosen to be noisy for CIFAR-10 and CIFAR-100, respectively. In this experiment, these noisy sources are chosen to be $100\%$ noisy so that we can understand how our method performs on data containing highly noisy sources. Noise is introduced based on the description given in Section 4. The ResNet 20 model described above is trained on this data and tasked with predicting the image class using cross-entropy loss. The model is trained for $40$ epochs in both cases, with the SGD optimiser and a learning rate of $0.1$, a momentum of $0.9$, and a weight decay of $0.0001$, in batches of size $128$. When training with LAP, we use $H=50$, $\delta=1.0$, and $\lambda=1.0$ for CIFAR-10 and $H=25$, $\delta=0.5$, and $\lambda=1.0$ for CIFAR-100 chosen using the validation data (which are made noisy using the same procedure as the training data) after training on data with random label noise. When training on CIFAR-100 we found that using a warm-up of $100$ steps improved performance.
  • F-MNIST: All training data is randomly split into $10$ sources, with $6$ chosen to be $100\%$ noisy. The low capacity MLP model is trained for $40$ epochs using the Adam optimiser, and a learning rate of $0.001$ on batches of size $200$. When training with LAP, we use $H=25$, $\delta=0.5$, and $\lambda=0.5$ chosen using the validation data (which is made noisy using the same procedure as the training data) after training on data with random label noise.
In our implementation, each data point contained an observation, the generating data source, and a target. The features and labels were passed to the model for training, and the sources were used to calculate the weighting to apply to the loss at each data point.
To allow fairness in comparison with ARFL, global updates on this model were performed the same number of times as the number of epochs other models were trained for, with all clients being trained on data for a single epoch before each global update. All other parameters for the ARFL model are kept the same. When IDPA was tested, the parameters were chosen as in Wang et al. [16] using the default parameters in the implementation. When testing “Co-teaching”, the parameters are chosen as in Han et al. [15] where possible, and scaled proportionally by the change in the number of epochs between their setting and our setting where they depended on the total number of epochs. The forgetting rate was set as the default given in the implementation code, $0.2$, since we do not assume access to the true noise rate. For CDR, the parameters for CIFAR-10 and 100 were chosen as given in Xia et al. [14], and for F-MNIST, we used the same parameters as CIFAR-10. When testing Label Smoothing [18], we used a smoothing rate of $0.1$ for CIFAR-10, CIFAR-100, and F-MNIST as we found that using negative smoothing had adverse effects on accuracy.
Note that when training ARFL, since it is a federated learning method, each client is trained on sources separately.
@@
For all of the following experiments, to reduce the computational cost, we fixed the parameters of LAP to $H=25$, $\delta=1.0$, and $\lambda=0.8$. In practice, optimising these parameters should allow improved performance when using LAP over those presented in this work.
Results in Table 2. To produce the results presented here, all training data is randomly divided into $10$ sources, with $4$ chosen to be $100\%$ noisy. The high capacity CNN with contrastive learning presented in Li et al. [17] is trained for $25$ epochs, with all other parameters kept as in the original work. In addition to this model, a version (with the same parameters) is trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. These models were trained on batch sizes of $128$ using stochastic gradient descent with a learning rate of $0.02$, momentum of $0.9$, and weight decay $0.0005$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
@@
Results in Figure 5(a). This experiment allowed us to test the performance of using LAP on different numbers of sources with varied noise levels. Firstly, data is divided into $12$ sources based on the clinician performing the labelling of the ECG recording. Then, for a given number of sources (increasing along the $x$ axis), the noisy sources have noise levels that are set at linear intervals between $25\%$ and $100\%$. These sources are made noisy following the suggestions in Wong et al. [37] to simulate electromagnetic interference and using label flipping to simulate human error in labelling. Based on this data, we train the ResNet 1D model for $40$ epochs using the Adam optimiser with a learning rate of $0.001$ and a batch size of $64$. This model is trained with and without LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
@@
Results in Figure 5(b). Here, CIFAR-10 is used as the features and the labelling collected in Wei et al. [35] are used as the noisy labels to produce CIFAR-10N. The data is first randomly split into $10$ sources. Then, for a given number of sources (increasing along the $x$ axis), the noisy sources have noise levels that are set at linear intervals between $25\%$ and $100\%$, replacing the true CIFAR-10 labels with the real noisy labels from Wei et al. [35]. The high capacity CNN with contrastive learning presented in Li et al. [17] is trained for $25$ epochs, with all other parameters kept as in the original work. As before, a version of this model (with the same parameters) is trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. These models are trained on batch sizes of $128$ using stochastic gradient descent with a learning rate of $0.02$, momentum of $0.9$, and weight decay $0.0005$. Data is randomly assigned to mini-batches, with each batch containing multiple sources.
@@
Results in Table 3. In this set of experiments, we want to test the use of LAP in a natural language setting. Firstly, we load the training and testing sets and randomly split the training set into $10$ sources uniformly. We then choose $4$ of the sources to be $100\%$ unreliable and introduce noise through random labelling and random permuting the order of the text. We truncate or extend each review such that it contains exactly $256$ tokens. We use an LSTM to predict the sentiment of movie reviews by training the model for $40$ epochs with a batch size of $128$, a leaning rate of $0.001$, and the Adam optimiser [49]. A version of this model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources. Since IDPA, Co-teaching, CDR, and Label Smoothing are not bench-marked on this dataset, for the baseline-specific parameters, we used the same values as given for CIFAR-10.
@@
Results in Table 5. This experiment is designed to test the performance of LAP on a dataset with real-world data sources with a significant imbalance in both their size and class distributions (for further details in Appendix A.15) to better understand the robustness of our proposed method, since it calculates the log-likelihood of sources during training. We chose $30$ of the total $82$ raters to produce noisy labels. The Transformer Encoder was trained for $25$ epochs with a batch size of $256$ and a learning rate of $0.001$ using the Adam optimiser [49]. A version of this model (with the same parameters) is also trained using LAP with $H=25$, $\delta=1.0$, and $\lambda=0.8$. As before, data is randomly assigned to mini-batches, with each batch containing multiple sources. Since IDPA, Co-teaching, CDR, and Label Smoothing are not bench-marked on this dataset, for the baseline-specific parameters, we used the same values as given for CIFAR-10.